# MyFramelinecolor=black, outerlinewidth=2pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, backgroundcolor=gray!40

### v1

@@ 1 Function GetColor($g$):
Optimizing for Local Writes. A special case is that a server issues a write to an object that resides in its own heap partition. While the coherence protocol still guarantees safety, requiring moving an object in its local heap each time it is written clearly brings inefficiencies. To optimize for local writes, DistR adopts a pointer-coloring method, inspired by the design of concurrent garbage collectors in a managed runtime system such as JVM [1, 53]. Several utility pointer coloring functions are shown in Algorithm 3 which are used when dereferencing and dropping a reference. We reserve the first 16 bits of a global address as a “color”. The color value stored in the object’s owner gets incremented upon the expiration of a mutable reference, as detailed in Lines 11 in Algorithm 1. Any subsequent immutable borrow would look up the cache with the object’s global address. Even if the actual address remains the same, its color changes if a write has occurred. As such, the lookup would not return any stale copy from the local cache.
The 16-bit color field may overflow when the pointer keeps being borrowed for local writes on the same server. DistR implements a move-on-overflow strategy that moves the object to a new address and resets its color to zero once the maximum color value is reached ($2^{16}$), thereby preventing overflow and maintaining system integrity and performance.
Writing Unsafe Code in DistR. Rust allows developers to bypass compiler safety checks and write unsafe code for low-level operations such as accessing raw pointers and mutating shared variables at their own risk [45, 70]. Since DistR relies on SWMR semantics enforced by Rust’s ownership types, DistR ensures consistency and memory safety only in the “safe” Rust code. DistR does not cache objects in unsafe code but allows developers to implement their own cache. Developers must ensure that they do not violate consistency in unsafe code blocks where type safety is not enforced. This caution mirrors practices in other managed languages, like native code in Java and unsafe code in C#. To assist developers, DistR offers primitives such as dalloc, dread, and dwrite for managing data on the global DSM heap.
@@ 4.1.2 Adapting Rust Standard Libraries
To further reduce the barrier for programs to run distributively, we reimplement several standard Rust libraries atop DistR’s core memory constructs covering four categories: threading for distributed computation (std::thread), inter-thread channel for communication (std::sync::mpsc), reference-counted pointers for ownership sharing (std::sync::Rc and std::sync::Arc), and shared-state locks for concurrency control (std::sync::Mutex and std::sync::atomic).
@@
Threading. DistR’s threading library enables Rust threads to run distributively with two major adaptations. First, it enables distributed thread launching by re-implementing the spawn interface. Internally, it captures the thread body as a closure during compile time and forwards it to the runtime. During execution, the runtime launches the thread according to each server’s load (details in §4.2.1). Second, DistR performs implicit ownership transfers between the parent and the child threads at the start or the end of the child thread execution. Thanks to the distributed ownership transfer support provided by DistR’s memory model, the implementation in the threading library is hidden from developers and preserves type soundness and memory safety. Additionally, DistR is compatible with advanced thread utilities such as thread::scope, which allows for the spawning of scoped threads that can borrow non-static data. These utilities ensure that all threads are joined at the end of their scope and can internally utilize DistR’s functions for spawning and joining threads, thus extending their applicability to the distributed setting.
@@
Inter-Thread Channel. DistR extends Rust’s channel to connect two distributed threads for message passing. DistR internally builds a network-based message queue for cross-server messages. Benefiting from the shared global heap, Box pointers and references can be safely copied and remain valid across servers. Therefore, the sender can push an object into the channel as is without serialization, even if it may contain Box pointers. DistR forwards the object binary bytes to the receiver over the network, and the receiver can recover the object from the binary by direct type conversion without deserialization.
@@
Ownership Sharing. Rust allows multiple owners to share an object via reference-counted smart pointers, which count the number of live owners. In this case, smart pointers only have read access, and the object lifetime terminates when all owners die and the reference count hits zero. DistR does not require special treatment for Rc as it only allows ownership sharing inside a single thread. For Arc which shares ownership among multiple threads, DistR handles it in a similar way to immutable references with on-demand local caching and lazy eviction.
@@
Shared-State Concurrency. Rust supports shared-state concurrency, primarily through its atomics and mutexes, where threads commonly share an atomic-typed value or one mutex via ownership sharing (i.e., Arc). Unfortunately, the ownership model cannot type check concurrent read/write to shared states. Hence, Rust relies on an unsafe implementation in its standard library. §4.1.1 already provides a general discussion on writing unsafe code in DistR, and here we focus on DistR’s implementation for distributed shared states.
@@
Shared states create a unique challenge for DistR, as they may be replicated on multiple servers and those states must be synchronized among these servers. For example, an Arc<AtomicBool> may be replicated across different servers and used independently, causing multi-version issues if not synchronized properly. DistR addresses this inconsistency by allocating the actual value on the global heap and storing only the Box pointer in atomic types. This design allows atomics to be freely moved or replicated across servers while keeping a single version of the actual value. To synchronize concurrent operations on atomics, DistR adapts methods of atomic types to forward the operation as a message to the server storing the actual value, which serializes all operations with unsafe logic similar to the original Rust to guarantee atomicity and memory consistency. Similarly, DistR implements Mutex by allocating its metadata and owned object on the global heap and leaving only Box pointers in the mutex struct. Concurrent operations on mutexes are serialized on the server storing the mutex.
@@ 4.1.3 Affinity Annotations
To further improve performance, DistR allows developers to provide optional data affinity semantics via annotations. These annotations are useful for many datacenter applications that make extensive use of object-oriented data structures that require pointer-chasing to access. For instance, Memcached [55] uses a chained hash table where each hash bucket stores its KV pairs with a linked list. To find one KV pair from a bucket, Memcached has to iterate the linked list following the node pointers. However, frequent pointer chasing is unfavorable in a distributed setting, where each pointer dereference incurs additional runtime checks and potential cross-server traffic. It would be beneficial for the runtime to colocate them on the same server and schedule the computation there.
@@
Listing 3: A linked list implementation with TBox in DistR. The use of TBox ties list nodes one by one. Iterating a list will fetch all nodes together (if they are on another server), and henceforth accessing any node is guaranteed local.
Data-Affinity Pointer. To expose data affinity for clustered placement, DistR includes a new pointer type TBox for developers to “tie” a heap object to its owner. TBox shares the same interfaces as the ordinary Box and can be used as a drop-in replacement for Box. However, TBox enforces that the pointed-to object must reside on the same server as its owner. In other words, when its owner object is copied or moved, the object referenced by TBox will be copied or moved as well. TBox can be used in a nested manner to allow a large union of objects to be co-located. The DistR runtime fetches (i.e., copies or moves) them together in a single batch, leading to fewer network round-trips and higher throughput. TBox can also be assigned to and owned by a stack variable, in which case the referenced object is pinned onto the heap partition of the server that hosts the stack and cannot be moved. Dereferencing a TBox is thus guaranteed to be a local access—DistR optimizes it by skipping the runtime check.
Listing 3 presents a linked list implementation using TBox. Our linked list uses TBox (Line LABEL:lst:list:tbox-dec1) to specify the data affinity between consecutive nodes. As a result, all list nodes are chained with TBox, forming an affinity group. Line LABEL:lst:list:sum-sttLABEL:lst:list:sum-end define a sum function that calculates the total sum of all node values. Assuming the list is non-empty, Line 7 dereferences the pointer to the head node, and the DistR runtime checks the location of the head node and fetches the entire list of nodes together if they are not local. Next, accessing each node inside the loop body (Line 8LABEL:lst:list:loop-end) is guaranteed local and hence skips runtime checks. Compared to using Box directly, TBox makes both data fetching and accessing more efficient.
@@
Data-Affinity Thread. To expose the affinity between data and computation for thread scheduling, DistR extends its threading library with a spawn_to interface. spawn_to mirrors the ordinary spawn interface to spawn a new thread but takes an additional Box pointer argument, which indicates where the thread should be created. The runtime checks where the Box points to and creates the new thread on that same server. A common practice to use spawn_to is to pass the mostly-accessed object as the location indicator. Listing 4 presents how the distributed accumulator (shown in Listing 2) can use spawn_to to offload a thread to the same server as a.val resides. Line 5 hence performs local dereference to a.val inside a.add().
4.2 DistR Runtime System
DistR’s runtime system is the core component that manages memory and compute resources. It includes a runtime library (§4.2.1) that is linked to each application and launched on each server and a cluster-wise global controller (§4.2.2).
@@ 4.2.1 Application-Integrated Runtime
Communication Layer. The DistR runtime uses its communication layer to support the cache coherence protocol and cross-server memory accesses. The communication layer consists of a control plane and a data plane, both built with RDMA. The control plane leverages two-sided verbs to send and receive small control messages, and the receiver can perform the coherence logic upon receiving the message to minimize the end-to-end latency. The data plane, in contrast, is specialized for bulky data transfer with one-sided verbs. It fetches an object as a whole with a single RDMA message upon pointer dereferencing without interrupting the target server, minimizing both latency and CPU usage.
Heap Allocator. The DistR runtime provides standard memory allocation interfaces and always returns global addresses to the upper-level language abstractions. It prioritizes local memory allocation as long as the local heap partition has sufficient space. This strategy improves data locality by colocating an object with its creating thread.
When the local heap partition is depleted, DistR queries the global controller and allocates memory on the most vacant server. For remote memory allocation, it forwards the request to the target server by sending a message through the communication layer and returns the allocated address to the user. Memory deallocation follows a similar logic but it bypasses the controller and finds the server directly via the object’s global address. The allocator does not reserve separate space for the local cache. Instead, it manages the cache as part of the local heap partition and always allocates cached entries in the local heap partition. Under memory pressure, the allocator will scan the local cache and evict entries that are no longer used (i.e., reference count hits zero).
Thread Scheduler. The DistR thread scheduler runs in the user space and schedules threads locally to maximize CPU utilization. It also provides thread migration functionalities, facilitating the global controller to balance load between busy and vacant servers.
@@
The scheduler adopts the method of cooperative multitasking and context switches between threads non-preemptively. A running thread yields its control flow proactively when developers call await or reactively upon long-latency operations. Similarly to other systems [60, 65, 82], our scheduler handles context switches as function calls, which allow DistR to save only a few registers per thread.
The scheduler supports creating/migrating a thread to another server as well. To migrate a thread, DistR sends its function pointer, the saved register state, and its stack to the target server. Because each thread reserves its stack address range globally, DistR can copy the stack across servers without changing its address. Therefore, the thread can be easily resumed by directly calling the function pointer with the saved register state on the target server. DistR generates code for state transmission during the compile time for the scheduler to call upon thread migration.
@@ 4.2.2 Global Controller
The controller runs as a daemon process on the machine where the program is launched. It manages cluster resources and coordinates memory allocation and thread migration. It periodically pings each server to probe and record its resource usage (CPU and memory). It controls resource allocation in cooperation with the DistR runtime on each server. When allocating memory or creating a thread, the runtime will first query the controller, which chooses a target server following adaptive policies (discussed later), and then coordinate with the runtime on the target server to perform the actual operation. The controller also maintains a global table to track the location of each thread; the table is queried and updated by the scheduler when migrating a thread.
During program execution, servers may run into imbalanced loads when objects get relocated or new threads are created. DistR balances the load of each server by migrating threads from the busy server to less occupied ones, following an adaptive policy to minimize cross-server memory accesses. If a server is about to run out of memory (>90% memory usage), the controller keeps migrating the thread that consumes the most local heap memory until the pressure is resolved. If the server is under compute congestion (>90% CPU usage), the controller migrates threads that frequently access remote objects. The thread is then moved to the server it accesses the most unless the target server is also overloaded, in which case it moves to a vacant server instead.
@@ 4.2.3 Fault Tolerance
In DistR, the global heap can be replicated to tolerate failures. Replication creates copies for each heap partition at the same virtual address on backup servers. Threads, in contrast, are not replicated for efficiency and are only executed with the primary global heap. To maintain a synchronized view between the primary heap partition and its backup copy, a thread must additionally write back to the backup partition after each mutable borrow. However, our insight is that the thread can batch modifications to an object and delay the write-back until the object ownership is transferred to another server, which is the time point that the object becomes visible to threads on other servers. When a server with a primary heap partition fails, the controller will automatically promote its backup server to the primary and add a new backup server.
@@ 5 Implementation
The majority of DistR was implemented in Rust except for its communication library which is in C. We implemented DistR’s core language constructs as three Rust types (i.e., struct): Ref<T>, MutRef<T>, and DBox<T>. They serve as the counterpart for the original Rust &T, &mut T, and Box<T>, respectively. We implemented the coherence protocol with traits on these types, including Copy, Clone, and Drop, which are automatically embedded into the program source code and executed when references/pointers are created or destroyed. To support unmodified Rust programs, we changed the Rust compiler and added additional compilation passes to transform Rust references and Box pointers into corresponding types in DistR.
Our communication layer links libibverbs directly for fast and kernel-bypassing RDMA networking. We implemented a low-level C library that covers basic connection establishment and exposes high-level Rust interfaces for various RDMA verbs, including RDMA_READ, RDMA_WRITE, RDMA_SEND, RDMA_RECV, ATOMIC_FETCH_AND_ADD, and ATOMIC_CMP_AND_SWP. We primarily utilize one-sided READ and WRITE verbs for data transfers between servers, as they outperform the two-sided SEND/RECV counterparts—one-sided operations bypass the CPU and OS at the receiver side, whereas two-sided operations require the receiver to pre-post RECV verbs and await notification upon message arrival. For instance, when a remote object is accessed via mutable references, DistR copies the object to local memory using the READ verb. Upon dropping the reference, DistR updates the original owner Box to reflect the new address, a process executed using the WRITE verb. Conversely, two-sided SEND/RECV verbs are utilized for control message exchanges, such as establishing connections across servers. Atomic verbs ATOMIC_FETCH_AND_ADD, and ATOMIC_CMP_AND_SWP are primarily utilized for implementing shared states (e.g., atomic types and mutexes). DistR uses the RC (reliable connection) transport type to ensure reliable transmission and strict message ordering.

### v2

@@ 1 Function GetColor($g$):
Optimizing for Local Writes. A special case is that a server issues a write to an object that resides in its own heap partition. While the coherence protocol still guarantees safety, requiring moving an object in its local heap each time it is written clearly brings inefficiencies. To optimize for local writes, DRust adopts a pointer-coloring method, inspired by the design of concurrent garbage collectors in a managed runtime system such as JVM [1, 52]. Several utility pointer coloring functions are shown in Algorithm 3 which are used when dereferencing and dropping a reference. We reserve the first 16 bits of a global address as a “color”. The color value stored in the object’s owner gets incremented upon the expiration of a mutable reference, as detailed in Lines 11 in Algorithm 1. Any subsequent immutable borrow would look up the cache with the object’s global address. Even if the actual address remains the same, its color changes if a write has occurred. As such, the lookup would not return any stale copy from the local cache.
The 16-bit color field may overflow when the pointer keeps being borrowed for local writes on the same server. DRust implements a move-on-overflow strategy that moves the object to a new address and resets its color to zero once the maximum color value is reached ($2^{16}$), thereby preventing overflow and maintaining system integrity and performance.
Writing Unsafe Code in DRust. Rust allows developers to bypass compiler safety checks and write unsafe code for low-level operations such as accessing raw pointers and mutating shared variables at their own risk [44, 70]. Since DRust relies on SWMR semantics enforced by Rust’s ownership types, DRust ensures consistency and memory safety only in the “safe” Rust code. DRust does not cache objects in unsafe code but allows developers to implement their own cache. Developers must ensure that they do not violate consistency in unsafe code blocks where type safety is not enforced. This caution mirrors practices in other managed languages, like native code in Java and unsafe code in C#. To assist developers, DRust offers primitives such as dalloc, dread, and dwrite for managing data on the global DSM heap.
@@ 4.1.2 Adapting Rust Standard Libraries
To further reduce the barrier for programs to run distributively, we reimplement several standard Rust libraries atop DRust’s core memory constructs covering four categories: threading for distributed computation (std::thread), inter-thread channel for communication (std::sync::mpsc), reference-counted pointers for ownership sharing (std::sync::Rc and std::sync::Arc), and shared-state locks for concurrency control (std::sync::Mutex and std::sync::atomic).
@@
Threading. DRust’s threading library enables Rust threads to run distributively with two major adaptations. First, it enables distributed thread launching by re-implementing the spawn interface. Internally, it captures the thread body as a closure during compile time and forwards it to the runtime. During execution, the runtime launches the thread according to each server’s load (details in §4.2.1). Second, DRust performs implicit ownership transfers between the parent and the child threads at the start or the end of the child thread execution. Thanks to the distributed ownership transfer support provided by DRust’s memory model, the implementation in the threading library is hidden from developers and preserves type soundness and memory safety. Additionally, DRust is compatible with advanced thread utilities such as thread::scope, which allows for the spawning of scoped threads that can borrow non-static data. These utilities ensure that all threads are joined at the end of their scope and can internally utilize DRust’s functions for spawning and joining threads, thus extending their applicability to the distributed setting.
@@
Inter-Thread Channel. DRust extends Rust’s channel to connect two distributed threads for message passing. DRust internally builds a network-based message queue for cross-server messages. Benefiting from the shared global heap, Box pointers and references can be safely copied and remain valid across servers. Therefore, the sender can push an object into the channel as is without serialization, even if it may contain Box pointers. DRust forwards the object binary bytes to the receiver over the network, and the receiver can recover the object from the binary by direct type conversion without deserialization.
@@
Ownership Sharing. Rust allows multiple owners to share an object via reference-counted smart pointers, which count the number of live owners. In this case, smart pointers only have read access, and the object lifetime terminates when all owners die and the reference count hits zero. DRust does not require special treatment for Rc as it only allows ownership sharing inside a single thread. For Arc which shares ownership among multiple threads, DRust handles it in a similar way to immutable references with on-demand local caching and lazy eviction.
@@
Shared-State Concurrency. Rust supports shared-state concurrency, primarily through its atomics and mutexes, where threads commonly share an atomic-typed value or one mutex via ownership sharing (i.e., Arc). Unfortunately, the ownership model cannot type check concurrent read/write to shared states. Hence, Rust relies on an unsafe implementation in its standard library. §4.1.1 already provides a general discussion on writing unsafe code in DRust, and here we focus on DRust’s implementation for distributed shared states.
@@
Shared states create a unique challenge for DRust, as they may be replicated on multiple servers and those states must be synchronized among these servers. For example, an Arc<AtomicBool> may be replicated across different servers and used independently, causing multi-version issues if not synchronized properly. DRust addresses this inconsistency by allocating the actual value on the global heap and storing only the Box pointer in atomic types. This design allows atomics to be freely moved or replicated across servers while keeping a single version of the actual value. To synchronize concurrent operations on atomics, DRust adapts methods of atomic types to forward the operation as a message to the server storing the actual value, which serializes all operations with unsafe logic similar to the original Rust to guarantee atomicity and memory consistency. Similarly, DRust implements Mutex by allocating its metadata and owned object on the global heap and leaving only Box pointers in the mutex struct. Concurrent operations on mutexes are serialized on the server storing the mutex.
@@ 4.1.3 Affinity Annotations
To further improve performance, DRust allows developers to provide optional data affinity semantics via annotations. These annotations are useful for many datacenter applications that make extensive use of object-oriented data structures that require pointer-chasing to access. For instance, Memcached [55] uses a chained hash table where each hash bucket stores its KV pairs with a linked list. To find one KV pair from a bucket, Memcached has to iterate the linked list following the node pointers. However, frequent pointer chasing is unfavorable in a distributed setting, where each pointer dereference incurs additional runtime checks and potential cross-server traffic. It would be beneficial for the runtime to colocate them on the same server and schedule the computation there.
@@
Listing 3: A linked list implementation with TBox in DRust. The use of TBox ties list nodes one by one. Iterating a list will fetch all nodes together (if they are on another server), and henceforth accessing any node is guaranteed local.
Data-Affinity Pointer. To expose data affinity for clustered placement, DRust includes a new pointer type TBox for developers to “tie” a heap object to its owner. TBox shares the same interfaces as the ordinary Box and can be used as a drop-in replacement for Box. However, TBox enforces that the pointed-to object must reside on the same server as its owner. In other words, when its owner object is copied or moved, the object referenced by TBox will be copied or moved as well. TBox can be used in a nested manner to allow a large union of objects to be co-located. The DRust runtime fetches (i.e., copies or moves) them together in a single batch, leading to fewer network round-trips and higher throughput. TBox can also be assigned to and owned by a stack variable, in which case the referenced object is pinned onto the heap partition of the server that hosts the stack and cannot be moved. Dereferencing a TBox is thus guaranteed to be a local access—DRust optimizes it by skipping the runtime check.
Listing 3 presents a linked list implementation using TBox. Our linked list uses TBox (Line LABEL:lst:list:tbox-dec1) to specify the data affinity between consecutive nodes. As a result, all list nodes are chained with TBox, forming an affinity group. Line LABEL:lst:list:sum-sttLABEL:lst:list:sum-end define a sum function that calculates the total sum of all node values. Assuming the list is non-empty, Line 7 dereferences the pointer to the head node, and the DRust runtime checks the location of the head node and fetches the entire list of nodes together if they are not local. Next, accessing each node inside the loop body (Line 8LABEL:lst:list:loop-end) is guaranteed local and hence skips runtime checks. Compared to using Box directly, TBox makes both data fetching and accessing more efficient.
@@
Data-Affinity Thread. To expose the affinity between data and computation for thread scheduling, DRust extends its threading library with a spawn_to interface. spawn_to mirrors the ordinary spawn interface to spawn a new thread but takes an additional Box pointer argument, which indicates where the thread should be created. The runtime checks where the Box points to and creates the new thread on that same server. A common practice to use spawn_to is to pass the mostly-accessed object as the location indicator. Listing 4 presents how the distributed accumulator (shown in Listing 2) can use spawn_to to offload a thread to the same server as a.val resides. Line 5 hence performs local dereference to a.val inside a.add().
4.2 DRust Runtime System
DRust’s runtime system is the core component that manages memory and compute resources. It includes a runtime library (§4.2.1) that is linked to each application and launched on each server and a cluster-wise global controller (§4.2.2).
@@ 4.2.1 Application-Integrated Runtime
Communication Layer. The DRust runtime uses its communication layer to support the cache coherence protocol and cross-server memory accesses. The communication layer consists of a control plane and a data plane, both built with RDMA. The control plane leverages two-sided verbs to send and receive small control messages, and the receiver can perform the coherence logic upon receiving the message to minimize the end-to-end latency. The data plane, in contrast, is specialized for bulky data transfer with one-sided verbs. It fetches an object as a whole with a single RDMA message upon pointer dereferencing without interrupting the target server, minimizing both latency and CPU usage.
Heap Allocator. The DRust runtime provides standard memory allocation interfaces and always returns global addresses to the upper-level language abstractions. It prioritizes local memory allocation as long as the local heap partition has sufficient space. This strategy improves data locality by colocating an object with its creating thread.
When the local heap partition is depleted, DRust queries the global controller and allocates memory on the most vacant server. For remote memory allocation, it forwards the request to the target server by sending a message through the communication layer and returns the allocated address to the user. Memory deallocation follows a similar logic but it bypasses the controller and finds the server directly via the object’s global address. The allocator does not reserve separate space for the local cache. Instead, it manages the cache as part of the local heap partition and always allocates cached entries in the local heap partition. Under memory pressure, the allocator will scan the local cache and evict entries that are no longer used (i.e., reference count hits zero).
Thread Scheduler. The DRust thread scheduler runs in the user space and schedules threads locally to maximize CPU utilization. It also provides thread migration functionalities, facilitating the global controller to balance load between busy and vacant servers.
@@
The scheduler adopts the method of cooperative multitasking and context switches between threads non-preemptively. A running thread yields its control flow proactively when developers call await or reactively upon long-latency operations. Similarly to other systems [60, 65, 82], our scheduler handles context switches as function calls, which allow DRust to save only a few registers per thread.
The scheduler supports creating/migrating a thread to another server as well. To migrate a thread, DRust sends its function pointer, the saved register state, and its stack to the target server. Because each thread reserves its stack address range globally, DRust can copy the stack across servers without changing its address. Therefore, the thread can be easily resumed by directly calling the function pointer with the saved register state on the target server. DRust generates code for state transmission during the compile time for the scheduler to call upon thread migration.
@@ 4.2.2 Global Controller
The controller runs as a daemon process on the machine where the program is launched. It manages cluster resources and coordinates memory allocation and thread migration. It periodically pings each server to probe and record its resource usage (CPU and memory). It controls resource allocation in cooperation with the DRust runtime on each server. When allocating memory or creating a thread, the runtime will first query the controller, which chooses a target server following adaptive policies (discussed later), and then coordinate with the runtime on the target server to perform the actual operation. The controller also maintains a global table to track the location of each thread; the table is queried and updated by the scheduler when migrating a thread.
During program execution, servers may run into imbalanced loads when objects get relocated or new threads are created. DRust balances the load of each server by migrating threads from the busy server to less occupied ones, following an adaptive policy to minimize cross-server memory accesses. If a server is about to run out of memory (>90% memory usage), the controller keeps migrating the thread that consumes the most local heap memory until the pressure is resolved. If the server is under compute congestion (>90% CPU usage), the controller migrates threads that frequently access remote objects. The thread is then moved to the server it accesses the most unless the target server is also overloaded, in which case it moves to a vacant server instead.
@@ 4.2.3 Fault Tolerance
In DRust, the global heap can be replicated to tolerate failures. Replication creates copies for each heap partition at the same virtual address on backup servers. Threads, in contrast, are not replicated for efficiency and are only executed with the primary global heap. To maintain a synchronized view between the primary heap partition and its backup copy, a thread must additionally write back to the backup partition after each mutable borrow. However, our insight is that the thread can batch modifications to an object and delay the write-back until the object ownership is transferred to another server, which is the time point that the object becomes visible to threads on other servers. When a server with a primary heap partition fails, the controller will automatically promote its backup server to the primary and add a new backup server.
@@ 5 Implementation
The majority of DRust was implemented in Rust except for its communication library which is in C. We implemented DRust’s core language constructs as three Rust types (i.e., struct): Ref<T>, MutRef<T>, and DBox<T>. They serve as the counterpart for the original Rust &T, &mut T, and Box<T>, respectively. We implemented the coherence protocol with traits on these types, including Copy, Clone, and Drop, which are automatically embedded into the program source code and executed when references/pointers are created or destroyed. To support unmodified Rust programs, we changed the Rust compiler and added additional compilation passes to transform Rust references and Box pointers into corresponding types in DRust.
Our communication layer links libibverbs directly for fast and kernel-bypassing RDMA networking. We implemented a low-level C library that covers basic connection establishment and exposes high-level Rust interfaces for various RDMA verbs, including RDMA_READ, RDMA_WRITE, RDMA_SEND, RDMA_RECV, ATOMIC_FETCH_AND_ADD, and ATOMIC_CMP_AND_SWP. We primarily utilize one-sided READ and WRITE verbs for data transfers between servers, as they outperform the two-sided SEND/RECV counterparts—one-sided operations bypass the CPU and OS at the receiver side, whereas two-sided operations require the receiver to pre-post RECV verbs and await notification upon message arrival. For instance, when a remote object is accessed via mutable references, DRust copies the object to local memory using the READ verb. Upon dropping the reference, DRust updates the original owner Box to reflect the new address, a process executed using the WRITE verb. Conversely, two-sided SEND/RECV verbs are utilized for control message exchanges, such as establishing connections across servers. Atomic verbs ATOMIC_FETCH_AND_ADD, and ATOMIC_CMP_AND_SWP are primarily utilized for implementing shared states (e.g., atomic types and mutexes). DRust uses the RC (reliable connection) transport type to ensure reliable transmission and strict message ordering.